$SPAD/src/input cycles1.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool cycles1.output
)set message test on
)set message auto off
)clear all
-- Input generated from CycleIndicatorsXmpPage
)expose EVALCYC
--S 1 of 46
complete 1
--R
--R
--R   (1)  (1)
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 1
--S 2 of 46
complete 2
--R
--R
--R        1       1   2
--R   (2)  - (2) + - (1 )
--R        2       2
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 2
--S 3 of 46
complete 3
--R
--R
--R        1       1         1   3
--R   (3)  - (3) + - (2 1) + - (1 )
--R        3       2         6
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 3
--S 4 of 46
complete 7

--R
--R
--R   (4)
--R     1       1          1          1     2     1         1            1     3
--R     - (7) + - (6 1) + -- (5 2) + -- (5 1 ) + -- (4 3) + - (4 2 1) + -- (4 1 )
--R     7       6         10         10          12         8           24
--R   +
--R      1   2      1     2     1       2     1     4     1   3      1   2 3
--R     -- (3 1) + -- (3 2 ) + -- (3 2 1 ) + -- (3 1 ) + -- (2 1) + -- (2 1 )
--R     18         24          12            72          48         48
--R   +
--R      1      5      1    7
--R     --- (2 1 ) + ---- (1 )
--R     240          5040
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 4
--S 5 of 46
elementary 7
--R
--R
--R   (5)
--R     1       1          1          1     2     1         1            1     3
--R     - (7) - - (6 1) - -- (5 2) + -- (5 1 ) - -- (4 3) + - (4 2 1) - -- (4 1 )
--R     7       6         10         10          12         8           24
--R   +
--R      1   2      1     2     1       2     1     4     1   3      1   2 3
--R     -- (3 1) + -- (3 2 ) - -- (3 2 1 ) + -- (3 1 ) - -- (2 1) + -- (2 1 )
--R     18         24          12            72          48         48
--R   +
--R        1      5      1    7
--R     - --- (2 1 ) + ---- (1 )
--R       240          5040
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 5
--S 6 of 46
alternating 7
--R
--R
--R   (6)
--R     2       1     2    1           1   2      1     2     1     4     1   2 3
--R     - (7) + - (5 1 ) + - (4 2 1) + - (3 1) + -- (3 2 ) + -- (3 1 ) + -- (2 1 )
--R     7       5          4           9         12          36          24
--R   +
--R       1    7
--R     ---- (1 )
--R     2520
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 6

--S 7 of 46
cyclic 7
--R
--R
--R        6       1   7
--R   (7)  - (7) + - (1 )
--R        7       7
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 7
--S 8 of 46
dihedral 7
--R
--R
--R        3       1   3      1   7
--R   (8)  - (7) + - (2 1) + -- (1 )
--R        7       2         14
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 8
--S 9 of 46
graphs 5
--R
--R
--R   (9)
--R   1           1   2    1   2     1   3     1   4 2     1   3 4     1    10
--R   - (6 3 1) + - (5 ) + - (4 2) + - (3 1) + - (2 1 ) + -- (2 1 ) + --- (1  )
--R   6           5        4         6         8          12          120
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 9
--S 10 of 46
cap(complete 2**2, complete 2*complete 1**2)
--R
--R
--R   (10)  4
--R                                                      Type: Fraction(Integer)
--E 10
--S 11 of 46
cap(elementary 2**2, complete 2*complete 1**2)
--R
--R
--R   (11)  2
--R                                                      Type: Fraction(Integer)
--E 11
--S 12 of 46
cap(complete 3*complete 2*complete 1,complete 2**2*complete 1**2)

--R
--R
--R   (12)  24
--R                                                      Type: Fraction(Integer)
--E 12
--S 13 of 46
cap(elementary 3*elementary 2*elementary 1,complete 2**2*complete 1**2)
--R
--R
--R   (13)  8
--R                                                      Type: Fraction(Integer)
--E 13
--S 14 of 46
cap(complete 3*complete 2*complete 1,elementary 2**2*elementary 1**2)
--R
--R
--R   (14)  8
--R                                                      Type: Fraction(Integer)
--E 14
--S 15 of 46
eval(cup(complete 3*complete 2*complete 1, _
     cup(complete 2**2*complete 1**2,complete 2**3)))
--R
--R
--R   (15)  1500
--R                                                      Type: Fraction(Integer)
--E 15
--S 16 of 46
square:=dihedral 4
--R
--R
--R         1       3   2    1     2    1   4
--R   (16)  - (4) + - (2 ) + - (2 1 ) + - (1 )
--R         4       8        4          8
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 16
--S 17 of 46
cap(complete 2**2,square)
--R
--R
--R   (17)  2
--R                                                      Type: Fraction(Integer)
--E 17
--S 18 of 46

cap(complete 3*complete 2**2,dihedral 7)
--R
--R
--R   (18)  18
--R                                                      Type: Fraction(Integer)
--E 18
--S 19 of 46
cap(graphs 5,complete 7*complete 3)
--R
--R
--R   (19)  4
--R                                                      Type: Fraction(Integer)
--E 19
--S 20 of 46
s(x) == powerSum(x)
--R
--R                                                                   Type: Void
--E 20
--S 21 of 46
cube:=(1/24)*(s 1**8+9*s 2**4 + 8*s 3**2*s 1**2+6*s 4**2)
--R
--R   Compiling function s with type PositiveInteger ->
--R      SymmetricPolynomial(Fraction(Integer))
--R
--R         1   2    1   2 2    3   4     1   8
--R   (21)  - (4 ) + - (3 1 ) + - (2 ) + -- (1 )
--R         4        3          8        24
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 21
--S 22 of 46
cap(complete 4**2,cube)
--R
--R
--R   (22)  7
--R                                                      Type: Fraction(Integer)
--E 22
--S 23 of 46
cap(complete 2**3*complete 1**2,wreath(elementary 4,elementary 2))
--R
--R
--R   (23)  7
--R                                                      Type: Fraction(Integer)
--E 23
--S 24 of 46

cap(complete 2**3*complete 1**2,wreath(elementary 4,complete 2))
--R
--R
--R   (24)  17
--R                                                      Type: Fraction(Integer)
--E 24
--S 25 of 46
cap(complete 2**3*complete 1**2,wreath(complete 4,elementary 2))
--R
--R
--R   (25)  10
--R                                                      Type: Fraction(Integer)
--E 25
--S 26 of 46
cap(complete 2**3*complete 1**2,wreath(complete 4,complete 2))
--R
--R
--R   (26)  23
--R                                                      Type: Fraction(Integer)
--E 26
--S 27 of 46
x: ULS(FRAC INT,'x,0) := 'x
--R
--R
--R   (27)  x
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 27
--S 28 of 46
ZeroOrOne: INT -> ULS(FRAC INT, 'x, 0)
--R
--R                                                                   Type: Void
--E 28
--S 29 of 46
Integers: INT -> ULS(FRAC INT, 'x, 0)
--R
--R                                                                   Type: Void
--E 29
--S 30 of 46
ZeroOrOne n == 1+x**n
--R
--R                                                                   Type: Void
--E 30
--S 31 of 46

ZeroOrOne 5
--R
--R   Compiling function ZeroOrOne with type Integer ->
--R      UnivariateLaurentSeries(Fraction(Integer),x,0)
--R
--R              5
--R   (31)  1 + x
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 31
--S 32 of 46
Integers n == 1/(1-x**n)
--R
--R                                                                   Type: Void
--E 32
--S 33 of 46
Integers 5
--R
--R   Compiling function Integers with type Integer ->
--R      UnivariateLaurentSeries(Fraction(Integer),x,0)
--R
--R              5    10      11
--R   (33)  1 + x  + x   + O(x  )
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 33
--S 34 of 46
eval(ZeroOrOne, graphs 5)
--R
--R
--R                   2     3     4     5     6     7     8    9    10      11
--R   (34)  1 + x + 2x  + 4x  + 6x  + 6x  + 6x  + 4x  + 2x  + x  + x   + O(x  )
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 34
--S 35 of 46
eval(ZeroOrOne,dihedral 8)
--R
--R
--R                   2     3     4     5     6    7    8
--R   (35)  1 + x + 4x  + 5x  + 8x  + 5x  + 4x  + x  + x
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 35
--S 36 of 46
eval(Integers,complete 4)
--R
--R
--R   (36)

--R             2     3     4     5     6      7      8      9      10      11
--R   1 + x + 2x  + 3x  + 5x  + 6x  + 9x  + 11x  + 15x  + 18x  + 23x   + O(x  )
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 36
--S 37 of 46
eval(Integers,elementary 4)
--R
--R
--R   (37)
--R      6    7     8     9     10     11     12      13      14      15      16
--R     x  + x  + 2x  + 3x  + 5x   + 6x   + 9x   + 11x   + 15x   + 18x   + 23x
--R   +
--R        17
--R     O(x  )
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 37
--S 38 of 46
eval(ZeroOrOne,cube)
--R
--R
--R                   2     3     4     5     6    7    8
--R   (38)  1 + x + 3x  + 3x  + 7x  + 3x  + 3x  + x  + x
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 38
--S 39 of 46
eval(Integers,cube)
--R
--R
--R   (39)
--R               2     3      4      5      6       7       8       9       10
--R     1 + x + 4x  + 7x  + 21x  + 37x  + 85x  + 151x  + 292x  + 490x  + 848x
--R   +
--R        11
--R     O(x  )
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 39
--S 40 of 46
eval(Integers,graphs 5)
--R
--R
--R   (40)
--R               2     3      4      5      6       7       8       9       10
--R     1 + x + 3x  + 7x  + 17x  + 35x  + 76x  + 149x  + 291x  + 539x  + 974x
--R   +
--R        11
--R     O(x  )

--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 40
--S 41 of 46
eval(ZeroOrOne ,graphs 15)
--R
--R
--R   (41)
--R               2     3      4      5      6       7       8        9        10
--R     1 + x + 2x  + 5x  + 11x  + 26x  + 68x  + 177x  + 496x  + 1471x  + 4583x
--R   +
--R        11
--R     O(x  )
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 41
--S 42 of 46
cap(dihedral 30,complete 7*complete 8*complete 5*complete 10)
--R
--R
--R   (42)  49958972383320
--R                                                      Type: Fraction(Integer)
--E 42
--S 43 of 46
sf3221:= SFunction [3,2,2,1]
--R
--R
--R   (43)
--R      1          1     2     1   2     1            1     4     1   2
--R     -- (6 2) - -- (6 1 ) - -- (4 ) + -- (4 3 1) + -- (4 1 ) - -- (3 2)
--R     12         12          16        12           24          36
--R   +
--R      1   2 2     1     2      1       3     1     5     1    4     1   3 2
--R     -- (3 1 ) - -- (3 2 1) - -- (3 2 1 ) - -- (3 1 ) - --- (2 ) + -- (2 1 )
--R     36          24           36            72          192        48
--R   +
--R      1   2 4     1      6     1    8
--R     -- (2 1 ) - --- (2 1 ) + --- (1 )
--R     96          144          576
--R                                 Type: SymmetricPolynomial(Fraction(Integer))
--E 43
--S 44 of 46
cap(sf3221,complete 2**4)
--R
--R
--R   (44)  3
--R                                                      Type: Fraction(Integer)
--E 44

--S 45 of 46
cap(sf3221, powerSum 1**8)
--R
--R
--R   (45)  70
--R                                                      Type: Fraction(Integer)
--E 45
--S 46 of 46
eval(Integers, sf3221)
--R
--R
--R   (46)
--R      9     10     11      12      13      14      15       16       17       18
--R     x  + 3x   + 7x   + 14x   + 27x   + 47x   + 79x   + 126x   + 196x   + 294x
--R   +
--R         19      20
--R     432x   + O(x  )
--R                         Type: UnivariateLaurentSeries(Fraction(Integer),x,0)
--E 46
)spool
)lisp (bye)

———-

References^{[1] nothing}

13